Mr. Dinkelspiel bought $32.56 worth of stamps. He bought 20 more 19-cent stamps than 50-cent stamps. He bought twice as many 32-cent stamps as 19-cent stamps. How many of each kind of stamp did he buy?
Answer by jorel555(1290)

Let s be the number of 19-cent stamps bought. Then 50-cent stamps equal s-20; and 32-cent stamps equal 2s. So:
19s+50(s-20)+64s=3256
19s+50s-1000+64s=3256
133s=4256
s=4256/133=32
In all, 32 of the 19-cent stamps were bought, 12 of the 50-cent stamps, and 64 of the 32-cent stamps!!!!!!!!!